# Analytical bounds for decoy-state quantum key distribution with discrete phase randomization

### v1

@@ Abstract
Index Terms:
Quantum communication, quantum key distribution, parameter estimation, quantum cryptography.
@@ I Introduction
Quantum key distribution (QKD) aims to establish a secure communication channel between two distant parties, conventionally referred to as Alice and Bob. The theoretical security of QKD protocols is based on quantum mechanics [lo1999unconditional, shor2000simple], as well as a set of assumptions on the behavior and characteristics of the physical devices involved [hoi-kwong2005decoy]. In real-world scenarios, these assumptions are not necessarily fully met due to a mismatch between the models assumed in the security analysis and that of the actual hardware, leading to potential vulnerabilities [abellan2014ultra, tang2014experimental, pathak2023phase]. Consequently, accurate and comprehensive modeling of devices used is essential for conducting realistic security analyses and ensuring the robustness of QKD implementations. In this work, we focus on the specific issue of phase randomization when weak coherent pulses (WCPs) are employed in QKD protocols. In particular, we look at implementations that rely on discrete phase randomization (DPR), as opposed to continuous phase randomization (CPR). While the latter is an easier case to analyze, it would be more practical to establish alignment between the theory and practice in the DPR case.
@@
In many QKD systems, WCPs are widely adopted as a substitute for ideal single-photon sources [nahar2023imperfect, yang2024high], which require more advanced technology. WCPs, typically generated by attenuated laser sources, offer a practical and scalable solution due to their ease of generation, cost-effectiveness, and compatibility with existing optical communications infrastructure. A critical aspect of using WCPs securely lies in the application of phase randomization [li2022improving]. By using CPR, the output state of the WCP-based encoder can be modeled as a statistical mixture of photon-number (Fock) states, with a Poisson distribution. This transformation is fundamental to the decoy-state method [ma2005practical, lo2006security], which often enables tight estimation of the single-photon components and protects against certain eavesdropping strategies, such as photon-number-splitting attacks [huttner1995quantum, wang2005beating]. In the absence of phase randomization, the coherent states retain their phase coherence, invalidating the assumptions underpinning standard security proofs for the decoy state QKD and rendering the system susceptible to sophisticated quantum attacks [tang2013source, curras2022security].
@@
Despite its theoretical advantages, achieving ideal CPR in practical QKD systems remains a challenging task. Passive phase randomization techniques that rely on turning the laser sources on and off to generate random global phase values have limitations on the pulse rate because of the transient time needed to stabilize the laser source. Relying on free-running lasers in conjunction with amplitude modulators could also result in correlated phase values for consecutive pulses [curras2023security], especially at high repetition rates, again not meeting the required conditions for CPR [abellan2014ultra]. On the other hand, active phase randomization schemes, which employ external modulators to impose random phases, presume a uniform distribution of phase values over $0,2\pi] [$chen2020multiple, tang2014experimental, wang2023finite, tian2025reference]. In practice, however, these schemes are constrained by the precision and stability of the modulation hardware, making it challenging to realize full CPR. This discrepancy between theoretical models and practical implementations may introduce subtle vulnerabilities in the security of QKD protocols.
@@
As a pragmatic alternative, DPR has been proposed [cao2015discrete]. In DPR-based implementations, the global phase of each pulse is selected from a finite, predefined set of discrete values. Although, in this approach, we have more control on the phase values used, hence can better ensure the match between theory and implementation, the source model becomes more complicated. Specifically, the resulting quantum channel no longer conforms to the standard photon-number channel model that underpins conventional decoy-state analysis [hoi-kwong2005decoy, ma2005practical, sun2013practical, li2022improving]. This deviation invalidates the assumptions used in traditional security proofs [hoi-kwong2005decoy], necessitating the development of new theoretical frameworks and security analyses tailored to DPR cases.
@@
The challenge of incorporating DPR into QKD protocols has been addressed for both BB84 [cao2015discrete] and measurement-device-independent (MDI) QKD [cao2020discrete] protocols. However, a significant limitation of existing approaches is related to their parameter estimation, which typically requires solving complex numerical optimization problems. The computational burden of these methods increases substantially with the number of discrete phase slices, making them less practical for real-time implementations, or when there are restrictions on computational power and memory, which is the case in many IoT scenarios [wang2023finite, nahar2023imperfect].
@@
The structure of this paper is as follows. In Section II, we introduce the theoretical framework, including the distinctions between CPR and DPR, and their implications for BB84 and MDI-QKD protocols. Building on the concepts from [cao2015discrete], we define the optimization problems for parameter estimation and explain the relevant constraints in Section III. Our analytical solutions are then presented in Section IV. In Section V, the numerical and analytical techniques are compared in certain simulation scenarios. Finally, we conclude our paper in Section VI.
@@ II System Description
In this section, we first explain how the source is modeled in the CPR and DPR cases. We then describe BB84 and MDI QKD protocols with DPR in the particular case of vacuum+weak decoy states with phase encoding in $X$ and $Y$ bases [ma2005practical, cao2015discrete, cao2020discrete].
@@ II-A Source description
where $|n\rangle$, $n=0,1,$..., represents the Fock state in the optical mode defined by the pulse. In the CPR case, we assume that $\theta$ is uniformly distributed in $0,2\pi)$. The state generated by the WCP source can then be expressed by the following density matrix
@@
which is diagonal in the Fock basis. This model, known as the photon-number channel, is at the core of many security proofs, particularly those involving WCPs with decoy states [ma2005practical, tupkary2025qkd]. However, realizing perfect phase randomization in practice is technologically demanding. To address this, DPR has been proposed as an alternative.
In DPR, the global phase is uniformly selected from a discrete set of $D$ values, $\theta_{j}=2\pi j/D$, where $j=0,\ldots,D-1$, [cao2015discrete]. The source output can then be modeled by the following density matrix
@@ II-B BB84 QKD with DPR
We study the BB84 with the vacuum+weak decoy state protocol using DPR. We consider the phase-encoding implementation of this protocol as shown in Fig. 1. The interferometer at the encoder will generate two pulses, denoted as reference (r) and signal (s), with a phase difference $\{0,\pi\}$ in the $X$ basis, and $\{\pi/2,3\pi/2\}$ in the $Y$ basis.
@@
In the case of BB84 with DPR, the input state to the encoder for any intensity $\alpha\in\{0,\nu,\mu\}$ is uniformly selected from the set of coherent states $\{|e^{i2\pi j/D}\sqrt{2\alpha}\rangle,j=0,...,D-1\}$. At the receiver, Bob uses a similar Mach-Zehnder interferometer to choose between the $X$ and $Y$ bases, and to measure the phase difference between the reference and signal pulses. After the exchange of the raw key, Alice and Bob follow the typical steps in a QKD protocol to sift and reconcile their keys, and extract a secret key using privacy amplification. We only consider the asymptotic scenario in this work, which means that the intensity $\mu$ is the main intensity to be sent and the probability of sending the weak decoy state or the vacuum state is negligibly small. We also use the efficient BB84 protocol [cao2015discrete], where the $X$ basis is primarily used for secret key generation.
Figure 1: Schematic for the phase encoding BB84 protocol with DPR. Alice prepares a weak coherent pulse with a randomly chosen discrete phase $\theta\in\{2\pi j/D,j=0,...,D-1\}$. The prepared state is then passed through a Mach-Zehnder interferometer, using which an additional phase shift $\phi_{a}\in\{0,\pi/2,\pi,3\pi/2\}$ is applied to the signal arm. The resulting pulses are transmitted through a quantum channel to Bob. Upon reception, Bob randomly selects a basis by choosing the relative phase $\phi_{b}\in\{0,\pi/2\}$ and does the measurement. They will keep the states when one of the two detectors clicks. In the case of double clicks, Bob randomly assigns a bit to his key. Based on the click pattern and their basis choices, Alice and Bob obtain their raw key.
@@
To model the outcome of the encoder in Fig. 1, we use the diagonal format of the phase randomized input state as shown in Eq. 3. For a fixed intensity $\mu$ and an encoded input state $|\lambda_{k}\rangle$, with k=0,,D1k=0,...,D-1, the encoder output state in $X$ and $Y$ bases for bits 0 and 11 are given by [cao2015discrete]
[…]
@@
Here, $|0_{x}\rangle_{k}$ and $|1_{x}\rangle_{k}$ ($|0_{y}\rangle_{k}$ and $|1_{y}\rangle_{k}$) are logical qubits in the $X$ ($Y$) basis for key bits 0 and $1$, respectively, and for simplicity, we have not shown the normalization coefficients.
@@
For security purposes, it is essential that Eve cannot distinguish states in $X$ and $Y$ bases. In other words, the density operators for different bases should have an identical expression. This requirement is quantified by the fidelity between the corresponding density matrices [cao2015discrete]:
$\begin{split}F_{k,\text{BB84}}=&\text{tr}\sqrt{\sqrt{\rho_{xk}}\rho_{yk}\sqrt{\rho_{xk}}},\end{split}$ (7)
@@
$\rho_{xk}$ $=|0_{x}\rangle_{kk}\langle 0_{x}|+|1_{x}\rangle_{kk}\langle 1_{x}|,$
$\rho_{yk}$ $=|0_{y}\rangle_{kk}\langle 0_{y}|+|1_{y}\rangle_{kk}\langle 1_{y}|.$ (8)
The numerical results in [cao2015discrete] show that for $k=0$ and $k=1$, the fidelity can get close to 1 when $D$ increases, indicating low basis dependence. This behavior is consistent with previous findings and supports the use of these components for key generation [lo2006security]. However, for $k>1$, the fidelity significantly decreases (e.g., around $0.5$ for $k=2$ for parameter values used in [cao2015discrete]), indicating stronger basis dependence. This shows such states do not necessarily contribute much to the secure key rate. Therefore, in this work, we restrict key generation to the components with $k=0$ and $k=1$, where the basis dependence is often sufficiently low under practical conditions.
@@ II-C MDI QKD with DPR
Figure 2: Schematic for the phase encoding MDI QKD protocol with DPR. Both Alice and Bob independently prepare weak coherent pulses with randomly chosen discrete phase values $\theta_{a},\theta_{b}\in\{2\pi j/D,j=0,...,D-1\}$. Each prepared state is passed through a Mach-Zehnder interferometer, after which an additional phase shift $\phi_{a}$ or $\phi_{b}$ is applied to the signal arm to encode the quantum information, where $\phi_{a},\phi_{b}\in\{0,\pi/2,\pi,3\pi/2\}$. The resulting optical pulses are then transmitted through quantum channels to an untrustworthy intermediary, Charlie, who performs an interference measurement on the incoming states. Charlie publicly announces the detection results. Based on this announcement and their own encoding choices, Alice and Bob can post-select correlated events to extract a shared key bit.
Fig. 2 shows the setup for the MDI QKD protocol considered in this paper. As in the case of BB84, we consider the phase encoding implementation of this protocol for the vacuum+weak decoy variant [ma2012alternative]. Alice and Bob each uses a BB84 encoder as explained in Section II-B, hence the output state from each encoder for a $|\lambda_{k}\rangle$ input is given by Eq. 6. At the midpoint of the quantum channel, a third party, commonly referred to as Charlie, performs an interference measurement on the incoming reference and signal pulses by using single-photon detectors. A successful detection event is registered when only one of the two detectors clicks for both pairs of reference and signal pulses. Charlie then publicly announces the detection outcomes. Alice and Bob will decide on the shared bit based on the pattern of clicks reported by Charlie [ma2012alternative]. The post-processing steps are followed as described in Section II-B.
To ensure that the bases are mutually unbiased the density operators in the $X$ and $Y$ bases must be identical. This requirement is quantified by the fidelity between the corresponding density matrices [cao2020discrete] as follows:
$\begin{split}F_{k,\text{MDI}}=\text{tr}\sqrt{\sqrt{\rho_{xk}}\rho_{yk}\sqrt{\rho_{xk}}},\end{split}$ (9)
@@
$\begin{split}&\rho_{xk}\\ =&(|0_{x}\rangle_{kk}\langle 0_{x}|+|1_{x}\rangle_{kk}\langle 1_{x}|)_{A}\otimes(|0_{x}\rangle_{kk}\langle 0_{x}|+|1_{x}\rangle_{kk}\langle 1_{x}|)_{B},\\ &\rho_{yk}\\ =&(|0_{y}\rangle_{kk}\langle 0_{y}|+|1_{y}\rangle_{kk}\langle 1_{y}|)_{A}\otimes(|0_{y}\rangle_{kk}\langle 0_{y}|+|1_{y}\rangle_{kk}\langle 1_{y}|)_{B}.\end{split}$ (10)
Again, in practical regimes of operation, it has been observed that the main contribution to the key rate is because of the terms with $k=0$ or $k=1$ [cao2020discrete]. Our key rate calculations, in the next section, are, therefore, restricted to such cases.
Box 1: Nonlinear optimization problem for DPR BB84 protocol with vacuum+weak decoy states Generic Formulation Minimize $R^{\text{DPR}}_{\text{BB84}}=\sum_{k=\{0,1\}}Q_{x,k}^{\mu}[1-h_{2}(e_{x,k}^{p,\mu})]-fQ_{x}^{\mu}h_{2}(E_{x}^{\mu}).$ subject to $Q_{\gamma}^{\alpha}=\sum_{k=0}^{D-1}p_{k}^{\alpha}Y_{\gamma,k}^{\alpha},Q_{\gamma}^{\alpha}E_{\gamma}^{\alpha}=\sum_{k=0}^{D-1}p_{k}^{\alpha}e_{\gamma,k}^{b,\alpha}Y_{\gamma,k}^{\alpha},$ $|Y_{\gamma,k}^{\alpha_{1}}-Y_{\gamma,k}^{\alpha_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}},|e_{\gamma,k}^{b,\alpha_{1}}Y_{\gamma,k}^{\alpha_{1}}-e_{\gamma,k}^{b,\alpha_{2}}Y_{\gamma,k}^{\alpha_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}},$ $e^{p,\mu}_{x,k}\leq e^{b,\mu}_{y,k}+4\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})(1-2e^{b,\mu}_{y,k})$ $\quad\quad+4(1-2\Delta^{\mu}_{k})\sqrt{\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})e^{b,\mu}_{y,k}(1-e^{b,\mu}_{y,k})},$ where $\alpha,\alpha_{1},\alpha_{2}\in\{0,\nu,\mu\}$, $\gamma=\{x,y\}$ represents $X$ or $Y$ basis, $p_{k}^{\mu}$ is given in Eq. 5, $\epsilon_{\alpha_{1}\alpha_{2}}$ comes from Eq. 13, and $\Delta_{k}^{\mu}$ is defined in Eq. 15. Approximate Formulation As above, with the additional constraint that $Y_{\gamma,k}^{\alpha}=0$, for k=3,4,,D1k=3,4,...,D-1.

### v2

@@ I Introduction
Quantum key distribution (QKD) aims to establish a secure communication channel between two distant parties, conventionally referred to as Alice and Bob. The theoretical security of QKD protocols is based on quantum mechanics [13, 24], as well as a set of assumptions on the behavior and characteristics of the physical devices involved [8]. In real-world scenarios, these assumptions are not necessarily fully met due to a mismatch between the models assumed in the security analysis and that of the actual hardware, leading to potential vulnerabilities [1, 28, 22]. Consequently, accurate and comprehensive modeling of devices used is essential for conducting realistic security analyses and ensuring the robustness of QKD implementations. In this work, we focus on the specific issue of phase randomization when weak coherent pulses (WCPs) are employed in QKD protocols. In particular, we look at implementations that rely on discrete phase randomization (DPR), as opposed to continuous phase randomization (CPR). While the latter is an easier case to analyze, it would be more practical to establish alignment between the theory and practice in the DPR case.
@@
In many QKD systems, WCPs are widely adopted as a substitute for ideal single-photon sources [20, 36], which require more advanced technology. WCPs, typically generated by attenuated laser sources, offer a practical and scalable solution due to their ease of generation, cost-effectiveness, and compatibility with existing optical communications infrastructure. A critical aspect of using WCPs securely lies in the application of phase randomization. By using CPR, the output state of the WCP-based encoder can be modeled as a statistical mixture of photon-number (Fock) states, with a Poisson distribution. This transformation is fundamental to the decoy-state method [18, 14], which often enables tight estimation of the single-photon components and protects against certain eavesdropping strategies, such as photon-number-splitting attacks [9, 34]. In the absence of phase randomization, the coherent states retain their phase coherence, invalidating the assumptions underpinning standard security proofs for the decoy state QKD and rendering the system susceptible to sophisticated quantum attacks [27, 4].
@@
Despite theoretical advantages of CPR, achieving ideal CPR in practical QKD systems remains a challenging task. Passive phase randomization techniques that rely on turning the laser sources on and off to generate random global phase values have limitations on the pulse rate because of the transient time needed to stabilize the laser source. Relying on free-running lasers in conjunction with amplitude modulators could also result in correlated phase values for consecutive pulses [35], especially at high repetition rates, again not meeting the required conditions for CPR [1]. For example, in twin-field QKD settings, active phase randomization is employed to ensure security against coherent attacks and to reach long distances [31, 32]. In practice, however, these schemes are constrained by the precision and stability of the modulation hardware, making it challenging to realize full CPR [5]. Such discrepancies between theoretical models and practical implementations can introduce subtle vulnerabilities in the security of QKD protocols.
@@
As a pragmatic alternative, DPR has been proposed [2]. In DPR-based implementations, the global phase of each pulse is selected from a finite, predefined set of discrete values. Although, in this approach, we have more control on the phase values used, hence can better ensure the match between theory and implementation, the source model becomes more complicated. Specifically, the resulting quantum channel no longer conforms to the standard photon-number channel model that underpins conventional decoy-state analysis [8, 18, 25, 12]. This deviation invalidates the assumptions used in traditional security proofs [8], necessitating the development of new theoretical frameworks and security analyses tailored to DPR cases.
@@
The challenge of incorporating DPR into QKD protocols has been addressed for both BB84 [2] and measurement-device-independent (MDI) QKD [3] protocols. However, a significant limitation of existing approaches is related to their parameter estimation, which typically requires solving complex numerical optimization problems. The computational burden of these methods increases substantially with the number of discrete phase slices, making them less practical for real-time implementations, or when there are restrictions on computational power and memory, which is the case in many IoT scenarios [30, 20].
@@
The structure of this paper is as follows. In Section II, we introduce the theoretical framework, including the distinctions between CPR and DPR, and their implications for BB84 and MDI-QKD protocols. Building on the concepts from [2], we define the optimization problems for parameter estimation and explain the relevant constraints in Section III. Our analytical solutions are then presented in Section IV. In Section V, the numerical and analytical techniques are compared in certain simulation scenarios. Finally, we conclude our paper in Section VI.
@@ II System Description
In this section, we first explain how the source is modeled in the CPR and DPR cases. We then describe BB84 and MDI QKD protocols with DPR in the particular case of vacuum+weak decoy states with time-bin encoding in $Z$ and $X$ bases [18, 2, 3].
@@ II-A Source description
where $|n\rangle$, $n=0,1,$..., represents the Fock state of $n$ photons in the optical mode defined by the pulse.
In the CPR case, we assume that $\theta$ is uniformly distributed in $0,2\pi)$. The state generated by the WCP source can then be expressed by the following density matrix
@@
which is diagonal in the Fock basis. This model, known as the photon-number channel, is at the core of many security proofs, particularly those involving WCPs with decoy states [18, 29]. However, realizing perfect phase randomization in practice is technologically demanding. To address this, DPR has been proposed as an alternative.
In DPR, the global phase is uniformly selected from a discrete set of $D$ values, $\theta_{j}=2\pi j/D$, where $j=0,\ldots,D-1$, [2]. The source output can then be modeled by the following density matrix
@@ II-B BB84 QKD with DPR
We study the BB84 with the vacuum+weak decoy state protocol using DPR. We consider the time-bin encoding implementation of this protocol as shown in Fig. 1. In this scheme, the encoder output consists two pulses, denoted as reference (r) and signal (s). In $Z$ basis, a discrete phase randomized coherent state is sent in the reference (signal) pulse, with vacuum state in the other pulse, to represent bit 1 (0). In $X$ basis, a phase difference $\{0,\pi\}$ is introduced between the signal and reference pulses.
Fig. 1 shows a schematic way to implement the above encoding and its corresponding decoding procedures. For each intensity $\alpha\in\{0,\nu,\mu\}$, representing the vacuum, weak decoy, or the main signal, the input pulse to the encoder is uniformly chosen from the set of coherent states $\{|e^{i2\pi j/D}\sqrt{\alpha}\rangle,j=0,...,D-1\}$. Encoding in $Z$ basis can be achieved by generating the input pulse in horizontal (vertical) polarization, and setting the phase modulator to zero. In $X$ basis, the input pulse to the encoder has diagonal polarization, therefore splitting at the polarizing beam splitter (PBS), and the phase modulator introduces a phase difference $\{0,\pi\}$ between the signal and reference pulses. At the receiver, Bob either measures the time of arrival (for $Z$ basis) or uses a Mach-Zehnder interferometer to measure the phase difference between the reference and signal pulses ($X$ basis).
@@
After the exchange of the raw key, Alice and Bob follow the typical steps in the vacuum+weak decoy state protocol to sift and reconcile their keys, and extract a secret key using privacy amplification. We only consider the asymptotic scenario in this work, which means that the intensity $\mu$ is the main intensity to be sent and the probability of sending the weak decoy state or the vacuum state is negligibly small. We also use the efficient BB84 protocol [23], where the $Z$ basis is primarily used for secret key generation.
Figure 1: Schematic diagram of the time-bin encoding BB84 protocol with DPR. Alice prepares a weak coherent pulse with a randomly chosen discrete phase $\theta_{a}\in\{2\pi j/D,j=0,...,D-1\}$. By choosing the polarization of the input pulse, we can use a Mach-Zehnder interferometer to encode the input in $Z$ and $X$ bases. The resulting pulses are transmitted through a quantum channel to Bob. Upon receiving the pulses, Bob randomly selects his measurement basis. For the $Z$ basis, he determines the logical bit by recording whether the detection event occurs in the early or late time bin. For the $X$ basis, Bob measures the phase difference between r and s pulses. PBS: Polarizing beam splitter; PM: Phase modulator; OS: Optical switch; BS: Beam splitter; HWP: Half-wave plate.
@@
To model the outcome of the encoder in Fig. 1, we use the diagonal format of the phase randomized input state as shown in Eq. 3. For a fixed intensity $\mu$ and an encoded input state $|\lambda_{k}\rangle$, with k=0,,D1k=0,...,D-1, the encoder output state in $Z$ and $X$ bases for bits 0 and $1$ are, respectively, given by
$\begin{split}|0_{z}\rangle_{k}&=\sum_{j=0}^{D-1}e^{-i2\pi kj/D}|\sqrt{\mu}e^{i2\pi j/D}\rangle_{s}|0\rangle_{r},\\ |1_{z}\rangle_{k}&=\sum_{j=0}^{D-1}e^{-i2\pi kj/D}|0\rangle_{s}|\sqrt{\mu}e^{i2\pi j/D}\rangle_{r},\\ |0_{x}\rangle_{k}&=\sum_{j=0}^{D-1}e^{-i2\pi kj/D}|\sqrt{\mu/2}e^{i2\pi j/D}\rangle_{s}|\sqrt{\mu/2}e^{i2\pi j/D}\rangle_{r},\\ |1_{x}\rangle_{k}&=\sum_{j=0}^{D-1}e^{-i2\pi kj/D}|\sqrt{\mu/2}e^{i(2\pi j/D+\pi)}\rangle_{s}|\sqrt{\mu/2}e^{i2\pi j/D}\rangle_{r}.\end{split}$ (6)
@@
Here, $|0_{z}\rangle_{k}$ and $|1_{z}\rangle_{k}$ ($|0_{x}\rangle_{k}$ and $|1_{x}\rangle_{k}$) are logical qubits in the $Z$ ($X$) basis for key bits 0 and $1$, respectively, and for simplicity, we have not shown the normalization coefficients.
@@
For security purposes, it is essential that Eve cannot distinguish states in $Z$ and $X$ bases. In other words, the density operators for different bases should have an identical expression. This requirement is quantified by the fidelity between the corresponding density matrices [2]:
$\begin{split}F_{k,\text{BB84}}=&\text{tr}\sqrt{\sqrt{\rho_{zk}}\rho_{xk}\sqrt{\rho_{zk}}},\end{split}$ (7)
@@
$\rho_{zk}$ $=|0_{z}\rangle_{kk}\langle 0_{z}|+|1_{z}\rangle_{kk}\langle 1_{z}|,$
$\rho_{xk}$ $=|0_{x}\rangle_{kk}\langle 0_{x}|+|1_{x}\rangle_{kk}\langle 1_{x}|.$ (8)
Following the numerical methods in [2], we can verify that for $k=0$ and $k=1$, the above fidelity can, respectively, get close to $1/\sqrt{2}$ and 1 when $D$ increases. This behavior is consistent with previous findings and supports the use of these components for key generation [14]. However, for $k>1$, the fidelity significantly decreases (e.g., around $0.35$ for $k=2$ for parameter values used in [2]), indicating stronger basis dependence. This implies that such states may not necessarily contribute much to the secure key rate. Therefore, in this work, we restrict key generation to the components with $k=0$ and $k=1$, where the basis dependence is often sufficiently low under practical conditions.
@@ II-C MDI QKD with DPR
Figure 2: Schematic diagram of the time-bin encoding MDI QKD protocol with DPR. Both Alice and Bob independently prepare weak coherent pulses with randomly chosen discrete phase values $\theta_{a},\theta_{b}\in\{2\pi j/D,j=0,...,D-1\}$ with respect to their local phase references. Each prepared state is passed through a BB84 encoder as in Fig. 1. The resulting optical pulses are then transmitted through quantum channels to an untrustworthy intermediary, Charlie, who performs an interference measurement on the incoming states. Charlie publicly announces the detection results. Based on this announcement and their own encoding choices, Alice and Bob can post-select correlated events to extract a shared key bit.
Fig. 2 shows the setup for the MDI QKD protocol considered in this paper. As in the case of BB84, we consider the time-bin encoding implementation of this protocol for the vacuum+weak decoy variant [19]. Alice and Bob each uses a BB84 encoder as explained in Section II-B, hence the output state from each encoder for a $|\lambda_{k}\rangle$ input is given by Eq. 6. At the midpoint of the quantum channel, a third party, commonly referred to as Charlie, performs an interference measurement on the incoming reference and signal pulses by using single-photon detectors. A successful detection event is registered when only one of the two detectors clicks for both pairs of reference and signal pulses. Charlie then publicly announces the detection outcomes. Alice and Bob will decide on the shared bit based on the pattern of clicks reported by Charlie [19]. The post-processing steps are followed as described in Section II-B.
There are two additional points to mention for the particular implementation in Fig. 2. First, note that there could be a random offset between the phase reference at Alice versus that of Bob. The phase values in Eq. 6 are therefore with respect to a local phase reference. Second, given that the scheme in Fig. 2 relies on interference, polarization maintenance needs to be employed to ensure maximum interference at the middle beam splitter. We assume that this is the case in our analysis.
To ensure that the bases are mutually unbiased the density operators in the $Z$ and $X$ bases must be identical. This requirement is quantified by the fidelity between the corresponding density matrices [3] as follows:
$\begin{split}F_{k,\text{MDI}}=\text{tr}\sqrt{\sqrt{\rho_{zk}}\rho_{xk}\sqrt{\rho_{zk}}},\end{split}$ (9)
@@
$\begin{split}&\rho_{zk}\\ =&(|0_{z}\rangle_{kk}\langle 0_{z}|+|1_{z}\rangle_{kk}\langle 1_{z}|)_{A}\otimes(|0_{z}\rangle_{kk}\langle 0_{z}|+|1_{z}\rangle_{kk}\langle 1_{z}|)_{B},\\ &\rho_{yk}\\ =&(|0_{x}\rangle_{kk}\langle 0_{x}|+|1_{x}\rangle_{kk}\langle 1_{x}|)_{A}\otimes(|0_{x}\rangle_{kk}\langle 0_{x}|+|1_{x}\rangle_{kk}\langle 1_{x}|)_{B}.\end{split}$ (10)
Box 1: Nonlinear optimization problem for DPR BB84 protocol with vacuum+weak decoy states Generic Formulation Minimize $R^{\text{DPR}}_{\text{BB84}}=\sum_{k=\{0,1\}}Q_{z,k}^{\mu}[1-h_{2}(e_{z,k}^{p,\mu})]-fQ_{z}^{\mu}h_{2}(E_{z}^{\mu}).$ subject to $Q_{\gamma}^{\alpha}=\sum_{k=0}^{D-1}p_{k}^{\alpha}Y_{\gamma,k}^{\alpha},$ $Q_{\gamma}^{\alpha}E_{\gamma}^{\alpha}=\sum_{k=0}^{D-1}p_{k}^{\alpha}e_{\gamma,k}^{b,\alpha}Y_{\gamma,k}^{\alpha},$ $|Y_{\gamma,k}^{\alpha_{1}}-Y_{\gamma,k}^{\alpha_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}},\quad{k=0,\ldots,D-1}$ $|e_{\gamma,k}^{b,\alpha_{1}}Y_{\gamma,k}^{\alpha_{1}}-e_{\gamma,k}^{b,\alpha_{2}}Y_{\gamma,k}^{\alpha_{2}}|\leq\epsilon_{\alpha_{1}\alpha_{2}},\quad{k=0,\ldots,D-1}$ $e^{p,\mu}_{z,k}\leq e^{b,\mu}_{x,k}+4\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})(1-2e^{b,\mu}_{x,k})$ $\quad+4(1-2\Delta^{\mu}_{k})\sqrt{\Delta^{\mu}_{k}(1-\Delta^{\mu}_{k})e^{b,\mu}_{x,k}(1-e^{b,\mu}_{x,k})},k=0,1$ where $\alpha,\alpha_{1},\alpha_{2}\in\{0,\nu,\mu\}$, $\gamma=\{z,x\}$ represents $Z$ or $X$ basis, $p_{k}^{\mu}$ is given in Eq. 5, $\epsilon_{\alpha_{1}\alpha_{2}}$ comes from Eq. 13, and $\Delta_{k}^{\mu}$ is defined in Eq. 15. Approximate Formulation As above, with the additional constraint that $Y_{\gamma,k}^{\alpha}=0$, for k=3,4,,D1k=3,4,...,D-1.